Table 13-9 Output Total Cost 니c 0 40 10 60 20 90 30 130 40 180 50 240
5. Refer to Table 13-9. What is average variable cost when output is 50 units? $3.60 b. $4.00 $4.40 d. $4.80 a. с.
Answer #1

Answer : The answer is option b.

Here when output is 0 unit then the total cost is $40. This means that here the fixed cost is $40.

When output is 50 units then total cost is $240.

At 50 units output level the total variable cost = Total cost - Fixed cost = 240 - 40 = $200. So,

At 50 units output level the average variable cost = Total variable cost / Output level = 200 / 50 = $4.

Therefore, option b is correct.